# Relative Pose Estimation for Nonholonomic Robot Formation with UWB-IO Measurements

### v2

@@ Abstract
This article studies the problem of distributed formation control for multiple robots by using onboard ultra wide band (UWB) distance and inertial odometer (IO) measurements. Although this problem has been widely studied, a fundamental limitation of most works is that they require each robot’s pose and sensor measurements are expressed in a common reference frame. However, it is inapplicable for nonholonomic robot formations due to the practical difficulty of aligning IO measurements of individual robot in a common frame. To address this problem, firstly, a concurrent-learning based estimator is firstly proposed to achieve relative localization between neighboring robots in a local frame. Different from most relative localization methods in a global frame, both relative position and orientation in a local frame are estimated with only UWB ranging and IO measurements. Secondly, to deal with information loss caused by directed communication topology, a cooperative localization algorithm is introduced to estimate the relative pose to the leader robot. Thirdly, based on the theoretical results on relative pose estimation, a distributed formation tracking controller is proposed for nonholonomic robots. Both 3D and 2D real-world experiments conducted on aerial robots and grounded robots are provided to demonstrate the effectiveness of the proposed method.
Index Terms:
Relative localization, nonholonomic robots, local reference frame, formation control.
@@ I Introduction
Distributed formation control for robot swarms has received significant attention in recent years due to its wide range of potential applications [1]. The objective of formation control is to guide robot swarms from an initial configuration to a specified formation through distributed interactions. To achieve this objective, various methods have been explored, including those based on consensus [2], artificial light field [3] and mean-shift exploration [4]. However, most of these methods rely on external localization systems, such as GPS, which are not well-suited for practical applications. Moreover, external localization systems may be unavailable in certain scenarios, such as indoor or underground environments. In summary, compared to relying on external localization systems, robots that depend solely on onboard measurements are more autonomous.
@@
Extensive researches adopted sensor measurements to estimate relative poses between neighboring robots. Based on the type of sensor utilized, formation control methods relying on local measurements can be roughly classified into two categories: relative position, (bearing or angle-based) methods, and distance-based methods. For the first class methods, depth camera or LADAR can be used to achieve ralative measurements between neighboring robots. Distributed formation control task can be accomplished when the inter-robot measurement topology satisfies specific conditions such as bearing rigidity [5] or non-collinear neighors condition [6, 7]. To relax the topology condition, recent works [8, 9] enhance the observability of relative pose of neighbor robots by more effectively utilizing visual measurement information. However, all these methods require expensive and complex sensors and object detection algorithms, significantly increasing both complexity and hardware costs. Additionally, these methods face challenges like limited field of view and high environmental requirements, including lighting conditions [10].
@@
Figure 1: Geometric relationship between the displacement and UWB distance measurements of the two robots. A. All the robot have a global orientation [11, 12, 13, 14, 15, 16, 17]. B. Each robot is in its local frame.
@@
Distance measurement sensors like Ultra Wide Band (UWB) inherently address the challenges [18, 19]. One of the primary difficulties of formation control that relies on distance measurements is accurately estimating. To estimate relative position, traditional methods typically derive the leader’s relative position using distance measurements from multiple neighbors [20]. However, this method inevitably imposes conditional constraints on the measurement topology, such as the requirement of rigidity. Similarly, a common strategy is to place multiple UWB tags on each robot [21, 22, 23] or install UWB anchors in the environment [24], which may sacrifice the scalability of the robot swarm. Another strategy involves introducing the inertial odometer (IO) measurement to achieve relative localization between robots [11]. Using a relative position estimator based on UWB distance and local IO measurements, distributed formation control of robot swarms has been achieved [12, 13, 14, 25]. To handle the sensor characteristics, the IO scale factor is estimated in [15], and an abnormal data removal mechanism is proposed in [16].
@@
All the aforementioned methods address the inter-robot relative localization problem within a global reference frame, as illustrated in Fig. 1 A. However, for nonholonomic robots, relative pose estimation between robots must be conducted in a distributed manner within the local coordinate frame, as shown in Fig. 1 B. This poses significant challenges when relying solely on relative distance measurements. Fortunately, by increasing the IO displacement of each robot within its odometer frame, the relative yaw angle between the two robots can be estimated [26, 27, 10]. The core concept involves formulating an optimization problem that relates real-time distance measurements and the relative pose between robots. In these methods, only UWB distance and local IO measurements are adopted to estimate relative pose estimation between two robots. Nevertheless, all these methods [12, 14, 27, 10] require that the relative motion between robots satisfy the persistently excited (PE) condition to ensure the convergence of the relative localization estimators. In other words, these methods require that robots maintain a non-zero relative velocity at all times. This is typically achieved through the introduction of continuous excitation noise [14] and periodic oscillation signals [11, 10] of varying amplitudes into the control inputs of the robots. The disadvantage lies in the contradiction between the motion needed for PE localization enhancement and relatively stationary formation control tasks, which can result in uneven motion trajectories and diminish control performance. To address this issue, a rotating UWB tag has been used to achieve the PE condition [17]. However, this solution significantly increases the complexity of the sensor equipment on each robot and raises the failure rate. So far, there are few studies focused on the implementation of distributed formation control for multiple nonholonomic robots using only UWB and IO measurements.
@@
The primary main challenge lies in achieving distributed cooperative pose estimation based on relative distance measurements without necessitating the PE condition for robot relative motion. This article aims to address the above challenge, and the contributions can be summarized as follows. First, the proposed distributed formation control law relies solely on inter-robot distance measurements. Unlike existing formation control methods for nonholonomic robots that require knowledge of inter-robot relative positions [1, 7, 28] or relative orientations [7, 29], our method utilizes only UWB distance and IO measurements. This reduces the sensor requirements for each robot and broadens applicability across various scenarios. Second, to facilitate relative pose estimation between robots in a local frame, a concurrent-learning based relative pose estimator is proposed. Different from recent methods[12, 14, 1] that estimate relative position or orientation in a global frame, our approach focuses on estimating relative pose in a local context. In contrast to PE relative motion based methods [14, 12, 27], our method effectively incorporates both current and historical information to relax the PE condition to a sufficiently excited (SE) condition in the relative pose estimation problem. This can effectively improve the smoothness of robot trajectories and enhance control performance. Additionally, the proposed data-driven methods demonstrate superior noise robustness compared to those relying solely on current information [14]. We provide both theoretical analysis and real-world experiments involving mobile robots and drones are provided to validate the effectiveness of the proposed methods.
@@
Notations: A transformation matrix in frame $\{A\}$ is denoted as $\textbf{T}^{A}:=\begin{bmatrix}\textbf{R}^{A},&\textbf{p}^{A}\\ \textbf{0}^{T},&1\\ \end{bmatrix}\in SE(3)$, where $\textbf{p}^{A}\in\mathcal{R}^{3}$ and $\textbf{R}^{A}\in SO(3)$ are the position vector and rotation matrix in frame $\{A\}$, respectively. Denote $\textbf{T}^{A}_{B}$ and $\textbf{R}^{A}_{B}$ as the transformation and rotation matrices from frame $\{B\}$ to frame $\{A\}$, respectively. For vector x, $x[i]$ and $x[i:j]$ denote the $i$-th element and the elements from $i$ to $j$ of x. For ease of reading, we list the main symbols of this article in Appendix A-A.
@@ II-A Robot Models
Consider a group of $N+1$ nonholonomic robots in $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$. For $i=0,...,N$, the kinematics of robot $i$ is described as
$\dot{\textbf{p}}^{\mathcal{O}_{i}}_{i}(t)&=\begin{bmatrix}\cos\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t),&0\\ \sin\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t),&0\\ 0,&1\end{bmatrix}\begin{bmatrix}v_{i,h}\\ v_{i,z}\end{bmatrix}\\ \dot{\phi}_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)&=w_{i}.\end{aligned}\right.$ (1)
@@ II-B UWB Measurement and Communication Topology
Each robot $i$ is equipped with a UWB module to measure the real-time distance and communicate with each other. Define the position of robot $j$ in frame $\mathcal{O}_{i}$ at time $t$ as $\textbf{p}^{\mathcal{O}_{i}}_{j}(t)$. Then the relative position between robot $i$ and robot $j$ in frame $\mathcal{O}_{i}$ can be denoted as $\textbf{p}^{\mathcal{O}_{i}}_{ij}(t)=\textbf{p}^{\mathcal{O}_{i}}_{i}(t)-\textbf{p}^{\mathcal{O}_{i}}_{j}(t)$. With the UWB module, each robot $i$ can measure the real-time distance $d_{ij}(t)\triangleq\lVert\textbf{p}^{\mathcal{O}_{i}}_{ij}(t)\rVert$ from itself to its neighbor $j$ and communicate with each other. The interaction network among the robots is described by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{0,1,\ldots,N\}$ and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Robot 00 is the leader robot. If $(i,j)\in\mathcal{E}$, robot $i$ and $j$ are neighbors, neighbors can communicate and measure distance with each other. The neighbor set of robot $i$ is denoted as $\mathcal{N}_{i}=\{j\in\mathcal{V}:(i,j)\in\mathcal{E}\}$.
@@ II-C Control Objective
$&\lim_{t\to\infty}\theta_{\Sigma_{i}}^{\Sigma_{0}}(t)=0\\ &\lim_{t\to\infty}\textbf{p}^{\mathcal{O}_{0}}_{i0}(t)-\textbf{p}_{i}^{*}=0,\end{aligned}\right.$ (2)
@@
where $\textbf{x}\in\mathbb{R}^{n_{x}}$ is the state and $\delta\in\mathbb{R}^{n_{\delta}}$ is an exogenous signal. $f(t,\textbf{x})$ and $g(t,\textbf{x},\bm{\delta})$ are piecewise continues in $t$ and locally Lipschitz in x and $(\textbf{x},\bm{\delta})$, respectively. Normal system $=f(t,\textbf{x})$ x ) in (3) is uniformly exponentially stable and $g(t,\textbf{x},\bm{\delta})$ satisfies $\lVert g(t,\textbf{x},\bm{\delta})\rVert\leq a_{1}\lVert\bm{\delta}\rVert\lVert\textbf{x}\rVert+a_{2}\lVert\bm{\delta}\rVert$, with two positive constants $a_{1}$ and $a_{2}$. Then system (3) uniformly asymptotically stable when $\lVert\bm{\delta}\rVert$ converges to 0 asymptotically.
@@ III-A Problem Formulation of Relative Localization
$-\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})^{T}\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})+d_{ij}(t_{k+1})^{2}\Big{)}+\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})^{T}\textbf{R}^{\mathcal{O}_{i}}\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})^{T}.$
@@
$\epsilon_{ij}(t_{k})=\hat{\mathbf{\Theta}}_{ij}(t_{k})\mathbf{\Phi}_{ij}(t_{k})-y_{ij}(t_{k})=\tilde{\mathbf{\Theta}}_{ij}(t_{k})\mathbf{\Phi}_{ij}(t_{k}).$ (12)
@@ III-B Relative Localization to Neighboring Robot
$\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t):=\mathrm{Norm}\left(\begin{bmatrix}\hat{\mathbf{\Theta}}_{ij}[6],&-\hat{\mathbf{\Theta}}_{ij}[7],&0\\ \hat{\mathbf{\Theta}}_{ij}[7],&\hat{\mathbf{\Theta}}_{ij}[6],&0\\ 0,&0,&1\end{bmatrix}\right),$ (14)
@@
Assumption 1 indicates that the data matrix $S_{ij}$ records sufficient information, which is also known as SE condition. Compared to the PE condition, SE condition is much weaker and can also be found in many data-driven estimation results like [30] and [31]. We also analyze the observability of relative pose when the relative motion between neighboring robots is different, which can be found in Appendix 15-D. The main results regarding the convergence of the estimator (13) are presented in the following theorem.
@@
In application, both UWB and odometer sensors have measurement errors. With the effective outlier detection algorithms, most measurement errors can be considered that the measurement error is bounded. According to Theorem 1 in [32], when the measurement error is bounded, there is a linear relationship between the estimation error and the upper bound of the measurement error.

### v4

@@ Abstract
This article studies the problem of distributed formation control for multiple robots by using onboard Ultra Wide Band (UWB) distance and inertial odometer (IO) measurements. Although this problem has been widely studied, a fundamental limitation of most works is that they require each robot’s pose and sensor measurements are expressed in a common reference frame. However, it is inapplicable for nonholonomic robot formations due to the practical difficulty of aligning IO measurements of individual robot in a common frame. To address this problem, firstly, a concurrent-learning based estimator is firstly proposed to achieve relative localization between neighboring robots in a local frame. Different from most relative localization methods in a global frame, both relative position and orientation in a local frame are estimated with only UWB ranging and IO measurements. Secondly, to deal with information loss caused by directed communication topology, a cooperative localization algorithm is introduced to estimate the relative pose to the leader robot. Thirdly, based on the theoretical results on relative pose estimation, a distributed formation tracking controller is proposed for nonholonomic robots. Both 2D and 3D real-world experiments conducted on aerial robots and grounded robots are provided to demonstrate the effectiveness of the proposed method.
@@ I Introduction
Distributed formation control for robot swarms has received significant attention in recent years due to its wide range of potential applications [1, 2, 3, 4]. The objective of formation control is to guide robot swarms from an initial configuration to a specified formation through distributed interactions. To achieve this objective, various methods have been explored, including those based on consensus [5], artificial light field [6] and mean-shift exploration [7]. However, most of these methods rely on external localization systems, such as GPS, which are not well-suited for practical applications. Moreover, external localization systems may be unavailable in certain scenarios, such as indoor or underground environments. In summary, compared to relying on external localization systems, robots that depend solely on onboard measurements are more autonomous.
@@
Extensive researches adopted sensor measurements to estimate relative poses between neighboring robots. Based on the type of sensor utilized, formation control methods relying on local measurements can be roughly classified into three categories: relative-position, relative-bearing (or angle), and relative-distance based methods. For the first class methods, depth camera or LADAR can be used to achieve ralative measurements between neighboring robots. Distributed formation control task can be accomplished when the inter-robot measurement topology satisfies specific conditions such as bearing rigidity [8] or non-collinear neighors condition [9]. To relax the topology condition, recent work [10] enhance the observability of relative pose of neighbor robots by more effectively utilizing visual measurement information. However, all these methods require expensive and complex sensors and object detection algorithms, significantly increasing both complexity and hardware costs. Additionally, these methods face challenges like limited field of view and high environmental requirements, including lighting conditions [11].
@@
Figure 1: Geometric relationship between the displacement and UWB distance measurements of the two robots. A. All the robot have a global orientation [12, 13, 14, 15, 16, 17]. B. Each robot is in its local frame.
@@
Distance measurement sensors like Ultra Wide Band (UWB) inherently address the challenges [18, 19]. One of the primary difficulties of formation control that relies on distance measurements is accurately estimating. To estimate relative position, traditional methods typically derive the leader’s relative position using distance measurements from multiple neighbors [20]. However, this method inevitably imposes conditional constraints on the measurement topology, such as the requirement of rigidity. Similarly, a common strategy is to place multiple UWB tags on each robot [21, 22] or install UWB anchors in the environment [23], which may sacrifice the scalability of the robot swarm. Another strategy involves introducing the inertial odometer (IO) measurement to achieve relative localization among robots [12]. Using a relative position estimator based on UWB distance and local IO measurements, distributed formation control of robot swarms has been achieved in [13, 16, 14, 24]. To handle the sensor characteristics, an abnormal data removal mechanism is proposed in [16].
@@
All the aforementioned methods assume that the IO frames of all robots are directionally aligned, as illustrated in Fig. 1 A. This assumption is usually achieved by equipping robots with sensors such as compasses, which are highly susceptible to interference from ambient magnetic field variations [25]. For a nonholonomic robot with its own local frame, the orientation of this frame changes dynamically as the robot moves, resulting in no consistently aligned direction. As a result, relative pose estimation between nonholonomic robots must be conducted within local frame, as shown in Fig. 1 B, which is challenging relying solely on relative distance measurements. Fortunately, by increasing the IO displacement of each robot within its odometer frame, the relative yaw angle between the two robots can be estimated [11]. The core concept involves formulating an optimization problem that relates real-time distance measurements and the relative pose between robots. Nevertheless, all these methods [14, 26, 11] require that the relative motion between robots satisfy the persistently excited (PE) condition to ensure the convergence of the relative localization estimators. In other words, these methods require that robots maintain a non-zero relative velocity at all times. This is typically achieved through the introduction of continuous excitation noise [14] and periodic oscillation signals [12, 11] of varying amplitudes into the control inputs of the robots. The disadvantage lies in the contradiction between the motion needed for PE localization enhancement and relatively stationary formation control tasks, which can result in uneven motion trajectories and diminish control performance. To address this issue, a rotating UWB tag has been used to achieve the PE condition [17]. However, this solution significantly increases the complexity of the sensor equipment on each robot and raises the failure rate. So far, there are few studies focused on the implementation of distributed formation control for multiple nonholonomic robots using only UWB and IO measurements.
@@
The primary challenge lies in achieving distributed pose estimation based on relative distance measurements across different local frames without necessitating the PE condition. This article aims to address the above challenge, and the contributions can be summarized as follows. First, the proposed distributed formation control law relies solely on inter-robot distance measurements. Unlike existing formation control methods for nonholonomic robots that require knowledge of inter-robot relative positions [2, 9, 27] or relative orientations [9, 28], our method utilizes only UWB distance and IO measurements. This reduces the sensor requirements for each robot and broadens applicability across various scenarios. Second, to facilitate relative pose estimation between robots in a local frame, a concurrent-learning based relative pose estimator is proposed. Different from recent methods[14, 2] that estimate relative position or orientation in a global frame, our approach focuses on estimating relative pose in a local context. In contrast to PE relative motion based methods [14, 26], our method effectively incorporates both current and historical information to relax the PE condition to a sufficiently excited (SE) condition in the relative pose estimation problem. This can effectively improve the smoothness of robot trajectories and enhance control performance. Additionally, the proposed data-driven methods demonstrate superior noise robustness compared to those relying solely on current information [14]. We provide both theoretical analysis and real-world experiments involving mobile robots and drones are provided to validate the effectiveness of the proposed methods.
@@
Notations: A transformation matrix in frame $\{A\}$ is denoted as $\textbf{T}^{A}:=\begin{bmatrix}\textbf{R}^{A},&\textbf{p}^{A}\\ \textbf{0}^{T},&1\\ \end{bmatrix}\in SE(3)$, where $\textbf{p}^{A}\in\mathcal{R}^{3}$ and $\textbf{R}^{A}\in SO(3)$ are the position vector and rotation matrix in frame $\{A\}$, respectively. Denote $\textbf{T}^{A}_{B}$ and $\textbf{R}^{A}_{B}$ as the transformation and rotation matrices from frame $\{B\}$ to frame $\{A\}$, respectively. For vector x, $x[i]$ and $x[i:j]$ denote the $i$-th element and the elements from $i$ to $j$ of x. For ease of reading, we list the main symbols of this article in Appendix B-A.
@@ II-A Robot Models
Consider a group of $N+1$ nonholonomic robots in $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$. For i=0,,Ni=0,...,N, the kinematics of robot $i$ is described as
$\left\{\begin{aligned} \dot{\textbf{p}}^{\mathcal{O}_{i}}_{i}(t)&=\begin{bmatrix}\cos\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t),&0\\ \sin\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t),&0\\ 0,&1\end{bmatrix}\begin{bmatrix}v_{i,h}\\ v_{i,z}\end{bmatrix}\\ \dot{\phi}_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)&=w_{i}.\end{aligned}\right$. (1)
@@ II-B UWB Measurement and Communication Topology
Each robot $i$ is equipped with a UWB module to measure the real-time distance and communicate with each other. Define the position of robot $j$ in frame $\mathcal{O}_{i}$ at time $t$ as $\textbf{p}^{\mathcal{O}_{i}}_{j}(t)$. Then the relative position between robot $i$ and robot $j$ in frame $\mathcal{O}_{i}$ can be denoted as $\textbf{p}^{\mathcal{O}_{i}}_{ij}(t)=\textbf{p}^{\mathcal{O}_{i}}_{i}(t)-\textbf{p}^{\mathcal{O}_{i}}_{j}(t)$. With the UWB module, each robot $i$ can measure the real-time distance $d_{ij}(t)\triangleq\lVert\textbf{p}^{\mathcal{O}_{i}}_{ij}(t)\rVert$ from itself to its neighbor $j$ and communicate with each other. The interaction network among the robots is described by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{0,1,\ldots,N\}$ and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Robot 0 is the leader robot. If $(i,j)\in\mathcal{E}$, robot $i$ and $j$ are neighbors, neighbors can communicate and measure distance with each other. The neighbor set of robot $i$ is denoted as $\mathcal{N}_{i}=\{j\in\mathcal{V}:(i,j)\in\mathcal{E}\}$.
@@ II-C Control Objective
$\left\{\begin{aligned} &\lim_{t\to\infty}\theta_{\Sigma_{i}}^{\Sigma_{0}}(t)=0\\ &\lim_{t\to\infty}\textbf{p}^{\mathcal{O}_{0}}_{i0}(t)-\textbf{p}_{i}^{*}=0,\end{aligned}\right$. (2)
@@
where $\textbf{x}\in\mathbb{R}^{n_{x}}$ is the state and $\delta\in\mathbb{R}^{n_{\delta}}$ is an exogenous signal. $f(t,\textbf{x})$ and $g(t,\textbf{x},\bm{\delta})$ are piecewise continues in $t$ and locally Lipschitz in x and $(\textbf{x},\bm{\delta})$, respectively. Normal system $\dot{\textbf{x}}=f(t,\textbf{x})$ in (3) is uniformly exponentially stable and $g(t,\textbf{x},\bm{\delta})$ satisfies $\lVert g(t,\textbf{x},\bm{\delta})\rVert\leq a_{1}\lVert\bm{\delta}\rVert\lVert\textbf{x}\rVert+a_{2}\lVert\bm{\delta}\rVert$, with two positive constants $a_{1}$ and $a_{2}$. Then system (3) uniformly asymptotically stable when $\lVert\bm{\delta}\rVert$ converges to 0 asymptotically.
@@ III-A Problem Formulation of Relative Localization
$-\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})^{T}\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})+d_{ij}(t_{k+1})^{2}\Big)+\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})^{T}\textbf{R}^{\mathcal{O}_{i}}\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k}).$
@@
$\epsilon_{ij}(t_{k})=\hat{\mathbf{\Theta}}_{ij}^{T}(t_{k})\mathbf{\Phi}_{ij}(t_{k})-y_{ij}(t_{k})=\tilde{\mathbf{\Theta}}_{ij}^{T}(t_{k})\mathbf{\Phi}_{ij}(t_{k}).$ (12)
@@ III-B Relative Localization to Neighboring Robot
$\left\{\begin{aligned} &\hat{\mathbf{\Theta}}_{ij}(t)=\hat{\mathbf{\Theta}}_{ij}(t_{k}),\quad t_{k}<t\leq t_{k+1}\\ &\hat{\mathbf{\Theta}}_{ij}(t_{k+1})=\hat{\mathbf{\Theta}}_{ij}(t_{k})-\eta\sum_{m=1}^{\varsigma}\mathbf{\Phi}_{ij}(t_{m})\epsilon_{ij}(t_{m})\\ &-\eta\mathbf{\Phi}_{ij}(t_{k})\epsilon_{ij}(t_{k}).\end{aligned}\right$. (13)
@@
Assumption 1 indicates that the data matrix $S_{ij}$ records sufficient information, which is also known as SE condition. Compared to the PE condition, SE condition is much weaker and can also be found in many data-driven estimation results like [29] and [30]. We also analyze the observability of relative pose when the relative motion between neighboring robots is different, which can be found in Appendix A-D. The main results regarding the convergence of the estimator (13) are presented in the following theorem.
@@
In application, both UWB and odometer sensors have measurement errors. With effective outlier detection algorithms, most measurement errors can be considered as bounded. According to Theorem 1 in [31], when the measurement error is bounded, there is a linear relationship between the estimation error and the upper bound of the measurement error.